1. Find the following sum : 6 + 8 +8 +10 + 12 +.......+100.
2. Find the nth term 7,49,343,2401,...
4,12,20,28.....
2,6,12,20

Answer by Alan3354(69443)

1. Find the following sum : 6 + 8 +8 +10 + 12 +.......+100.
Does it have 2 8's?
If it has one 8:
From 6 to 100 inclusive is 95 terms. 47 are odd, 48 are even.
--> 24 pairs of terms that total 106, 6+100, 8+98, 10+96. etc
Sum = 24*106 = 2544
---------
If 2 8's, add 8 = 2552
------------
------------
2. Find the nth term 7,49,343,2401,...
7^n
------
4,12,20,28.....
8n + 4, n = 0,1,2...
--------
2,6,12,20
I'm working this one. Repost it, someone else might help.